# Advancing Multimodal Large Language Models in Chart Question Answering with Visualization-Referenced Instruction Tuning

### v1

@@ 1 Introduction
Building MLLMs tailored for CQA necessitates high-quality training datasets and benchmarks. Recent research[21, 54, 42] in this realm primarily concentrates on scaling up training datasets encompassing charts, data tables, and QA pairs, using manual labeling and data synthesis techniques. These efforts have enhanced the performance of MLLMs in traditional CQA benchmarks[51, 55]. However, bottlenecks have emerged, posing challenges for further improving performance and adapting to real-world scenarios. Simply scaling up the training dataset without implementing quality control measures poses significant challenges in training efficiency and the feasibility of integrating these data into general MLLM training.
@@
Drawing inspiration from the results, we introduce a novel data engine (Sect. 5) aimed at generating instruction-enhanced CQA datasets. This engine comprises a data-filtering component (Sect. 5.1), utilizing a classifier with fine-grained chart features to reveal distributions and filter existing chart datasets. To mitigate the bias in the chart distributions and generate unavailable chart tasks, we further design a data generation component (Sect. 5.2) employing a chart space-guided data augmentation strategy to ensure the inclusivity of real-world possible charts. We further enrich reasoning-oriented QAs for the generated charts, contributing to a new CQA dataset and benchmark (Sect. 5.3) that features a wider variety of chart types and more QAs with effective visual instructions.
Existing MLLMs, mostly relying on CLIP encoders trained on natural images, are not optimally suited for visualization charts due to inherent differences. Recognizing the limitations, we develop a new MLLM (Sect. 6) that unfreezes the vision encoders in CLIP to better adapt to chart-specific features. Our MLLM is trained using the newly curated CQA dataset with more effective visual instructions. Additionally, we incorporate a mixture-of-resolution adaptation strategy[49] to enhance the fine-grained recognition capabilities of chart elements. Quantitative experiments (Sect. 7) demonstrate that even trained on a dataset with 80 times less CQA data, our model consistently outperforms state-of-the-art CQA models on established benchmarks.
@@
  • An MLLM that outperforms existing open-source CQA models on existing CQA benchmarks and comparable to the best commercial models on our proposed benchmark.
@@ 3 Related Work
The advent of MLLMs has shifted the paradigm, achieving breakthroughs in visual question answering[4]. Notably, the open-source generic model Qwen-VL[5] demonstrates superior performance over all specialized chart models in the ChartQA benchmark[51], especially those posed by humans as opposed to machine-generated questions. Despite these advancements, our extensive empirical study has uncovered limitations in the current MLLMs’ ability to handle real-world CQA tasks, especially those that fall outside of the training data distribution. Rectifying these limitations necessitates the consideration of the visualization reference model[9] when constructing training data, which elucidates the pratical mapping process from raw data to final graphical representations. Accordingly, this study contributes to enhancing the performance of MLLMs in CQA by integrating knowledge from the visualization reference process into training data generation and augmentation.
Enhancing Capabilities of MLLMs. The enhancement of MLLMs in specific scenarios, such as medicine images and text-dense images, can be categorized into two primary approaches: model-centric works that aim to improve the performance and efficiency of vision encoders or projectors; data-centric works that try to improve the model performance by boosting the number and quality of training data. In data-centric advancements, several studies employ powerful LLMs (e.g., GPT-4[3]) to generate various instruction-format VL tasks, like caption generation[45]. Another line of studies has explored converting classical VL task datasets (e.g., COCO[39]) into an instruction-following format with pre-defined templates. Within this context, to enhance chart comprehension, ChartLLaMA[21] finetunes LLaVA with 160K instruction data generated by GPT-4. Similarly, ChartAst[54] crawls a huge amount of tables from arXiv and then uses tables to generate charts for large-scale chart-to-table pre-training. ChartAst also generates QA pairs based on the tables they collected. Despite these efforts, the factors that contribute to efficient instruction data for chart understanding are still unclear.
@@
In the visualization field, real-world image datasets like Beagle[6], VisImage[16], Vis30K[12], and multi-view visualizations [14], together with practical QA benchmarks for visual literacy test[33, 57, 19], have been introduced. The challenge lies in converting them into high-quality instruction data due to sparse label annotations. Our research draws upon methodologies that utilize GPT to generate code-format charts and associated instruction data. Specifically, we aim to guide the data generation process with the well-defined chart-task space[33] to contribute a dataset encompassing the real-world spectrum of chart features and QA tasks, thereby improving current MLLMs’ chart understanding capability.
@@ 4.1.2 Impacts of Distribution Bias
To gain deeper insights into the underlying reasons for the issue, we examine the responses generated by MLLMs equipped with CoT. Figure 3 illustrates three typical cases, highlighting deficiencies in three categories: recognition errors, inference errors for numerical calculations, and inference errors regarding chart knowledge. Multiple factors contribute to these errors. First, errors often occur for chart types common in visual literacy but rare in ChartQA, such as normalized stack bar charts. Additionally, uncommon questions in ChartQA, such as accurately determining a range of data values, may lead MLLMs to struggle to identify the correct range.
@@ 4.2.2 Results and Analysis
Table 2 shows the results of the ablation experiment of baseline and the fine-tuned MLLMs on different question types in ChartQA-H test set. Overall, models fine-tuned with more training data (individual vs. combinations datasets) achieve higher accuracy. Specifically, the inclusion of human-generated ChartQA-H dataset substantially enhances model performance across all question types. In contrast, ChartQA-M dataset is less effective and mainly improves data retrieval and visual questions. This difference further underscores the limited impact of data retrieval questions for tackling CQA challenges and the critical role of diverse, reasoning-intensive questions over simple recognition questions. Moreover, Chart2Table serves as an accompanying effective instruction task if the data tables are available.
@@ 5 Data Engine
Total existing 801,686 7,659,563
Filtered dataset 100,000 326,314
@@ 5.1 Data filtering
Given that studies have shown common pretrained visual feature extractors (e.g., CLIP-Vit[58]) are not sensitive to fine-grained chart attributes[68], conventional filtering approaches that sample data in the pretrained feature space lead to inhomogeneity in these attributes. Additionally, most existing datasets lack detailed annotations beyond coarse-grained chart types (e.g., bar, line, and pie), posing challenges for stratified sampling. To mitigate this issue, we construct classifiers to learn those attributes in a supervised manner and then perform stratified sampling based on the labels predicted by the classifiers.
@@ 5.1.2 Image Sampling and Instruction Data Sampling
Figure 4 [Uncaptioned image] illustrates the image sampling process. We employ the CLIP-Vit[58] and a color extractor [1] to extract the overall feature and color feature of each image and then concatenate the two feature vectors to formulate a joint embedding space. Inspired by Bunny[23] and SemDeDup[2], we cluster images into $k$ clusters via $k$-means within the joint embedding space, aiming to group charts with similar features. To ensure chart attribute balancing, we incorporate stratified sampling within each cluster. Specifically, we create strata within each cluster according to predicted chart attributes and further perform sampling in each stratum. We identify duplicates by constructing an undirected graph, where edges connect image pairs with cosine similarity above a specified threshold $\mathcal{\varepsilon}$, indicating high feature similarity. We streamline the process by directly retaining only the image with the lowest cosine similarity to the stratum’s centroid from each set of semantic duplicates, thereby effectively reducing dataset size while preserving diversity. Finally, we manually adjust $\mathcal{\varepsilon}$ to obtain 100K charts, ensuring an appropriate training cost.
@@ 5.3 Visualization-referenced Dataset and Benchmark
Overview of Dataset. Our generated dataset comprises 11 chart types and 8 task categories, as outlined in visual literacy research [33, 19, 57]. Table 4 illustrates the chart-task space. The generated dataset includes 8,163 table-chart pairs and 32,280 chart-QA pairs. By integrating the generated dataset with the filtered dataset (shown in Table LABEL:tab:existing-datasets), we ultimately produce a dataset of 467K, which includes 108K table-chart pairs and 359K chart-QA pairs. In contrast to earlier datasets that relied on single sources or template-driven designs, our approach effectively combines our generated dataset with existing datasets derived from a wide range of real-world and synthetic sources. This integration features a diverse array of visual encodings, referencing real-world charts from multiple sources. We have made deliberate efforts to ensure an equitable distribution of chart types and QA pairs, with a particular emphasis on underrepresented chart types and question types in current datasets, such as range determination and distribution characterization, to provide a more comprehensive and balanced resource. Detailed chart and QA examples can be found in Supplementary Section S2.
@@ 6.1 Model Architecture
High-resolution input by mixing vision encoders with tokens compression. The employed two-layer MLP effectively connects the feature space of vision encoders and LLMs but results in visual tokens that positively correlate with the image resolution. For instance, CLIP-Vit-L-14 results in 5,329 tokens for a 1,022×1,022 resolution image, as each token corresponds to a 14×14 image patch[49], which is computationally expensive for MLLMs. The popular query-based strategy for implementing high-resolution input, QFormer[37], requires large-scale pretraining to achieve vision-language alignments, which is impractical for visualization scenarios as high-quality data is scarce. Therefore, we adopt a resolution-adaptation strategy[49] to improve the resolution while supporting training on normal scale data. The strategy embeds high-resolution features into the low-resolution features via adapters, thus reducing visual feature tokens. Specifically, following the settings of LLaVA-HR[49], we integrate CLIP-ViT-L[58] and CLIP-ConvNeXt[47] as a mixture of vision encoders and then mix their features with an adaptation strategy, thus maintaining control over the length of the visual token sequence. The resolutions of ViT and CNN are set to 448×448 and 1,024×1,024, respectively.
@@ 6.2 Training Settings
Unfreezing the vision encoders. Former works[45, 21, 42] choose to freeze the vision encoder in the whole training process as the pre-trained CLIP is already good at capturing features of natural images. Their MLLM training target is aligning the features extracted by CLIP to the LLM embedding space by tuning the projector and LLM. However, former research[68] has found that CLIP performs much worse in visualization images, as its pretraining corpus has a relatively small amount of charts with coarse annotation, leading to limited chart recognition ability without further tuning. Unfreezing CLIP’s parameters enables better adaptation to chart features and improves the overall performance of MLLM’s chart understanding for improved chart recognition ability.
Training data. We skip the pretraining process and directly leverage the initial projector weight to conduct instruction tuning. Our study aims to improve general MLLM’s chart comprehension ability. Therefore, our training data consists of two parts: the 665K original instruction tuning data of LLaVA-1.5 and the 467K chart-related data described in Sect. 4.
Hyperparameters settings. AdamW[30] is used as the optimizer, and the learning rate (LR) and batch size are set to 2e-5 and 128, respectively. The training epoch is set to 1. The LR scheduler is cosine decay, with a warmup ratio of 0.03. The training is running on 16×NVIDIA A800 for approximately 25 hours.
@@ 7.1 Chart Attributes Classification
Table 6: Results on our benchmarks.
@@
No Filtered Data 91.0 61.4 76.2 90.3
No Generation Data 92.7 63.7 78.2 91.2
Our model (467K CQA data) 93.5 64.9 79.2 91.8
@@ 7.2.1 Benchmarks
ChartQA. The dataset information, the relaxed accuracy metric, and the prompt for short answers have been illustrated in Sect. 4.1.2.
@@ 7.2.3 Results
Table 7 presents the results of our model’s performance against other models. It demonstrates that our model consistently outperforms the baseline across all tasks. Particularly, we surpass the current leading models, but with only over 80 times less data, showcasing our data filtering and generation effectiveness. Moreover, the ablation studies present the effectiveness of unfreezing vision encoders and mixing vision encoders for high-resolution.
@@ 8 Discussion
Visual encoder enhancement. Our research finds that unfreezing the vision encoders substantially enhances the chart recognition capabilities of MLLMs, showing original CLIP-Vit’s under-performance in chart images[68]. An intuitive alternative design is replacing CLIP-Vit with an encoder pre-trained specifically on chart images. For example, ChartInstruct-LLaMA [53] substitute the CLIP-Vit in LLaVA with the UniChart vision encoder[52]. However, the researchers did not observe model performance improvements compared to Unichart. This highlights the superiority of generic vision encoders, which learn robust image interpreting capabilities (e.g., localization) from millions of natural images. Moreover, understanding some real-world charts requires broad vision knowledge. For instance, infographics may incorporate natural images to depict certain chart elements[69] vividly. Borrowing LLaVA-Med’s lesson[35] in initializing CLIP-Vit, developing a visualization-domain CLIP with enhanced basic chart interpretation performance is a promising future work.
@@
Insights on applying MLLMs to complex reasoning visualization tasks. Our research finds that current MLLMs still face challenges in analytical tasks (e.g., finding anomalies and determining range). Recently, referential question-answering [72, 13] has been shown to benefit the comprehension of complex spatial relationships. It requires annotating bounding boxes and arrows in images and referring to these elements in questions. This task was not considered in our training data because of the lack of chart data with referential annotations. However, referential QAs are common in real-world visual analytics and potentially beneficial for tasks like finding anomalies. For example, we can use bounding boxes to label anomaly points or highlight a range of data elements, boosting MLLMs’ understanding of relevant tasks. We place the exploration of referential and other possible complex QA formats tailored for charts in future work. Furthermore, end-to-end MLLM outputs inherently pose uncertainty. Incorporating golden tables and code of charts with MLLM for interaction applications may be robust in complex visual analytics scenarios.

### v2

@@ 1 Introduction
Building MLLMs tailored for CQA necessitates high-quality training datasets and benchmarks. Recent research[24, 57, 45] in this field primarily focuses on scaling up training datasets that include charts, data tables, and QA pairs, employing manual labeling and data synthesis techniques. These efforts have enhanced the performance of MLLMs in traditional CQA benchmarks[54, 58]. However, bottlenecks have emerged, posing challenges for further improving performance and adapting to real-world scenarios. Simply scaling up the training dataset without implementing quality control measures poses significant challenges in training efficiency and the feasibility of integrating these data into general MLLM training.
@@
Drawing inspiration from the results, we introduce a novel data engine (Sect. 5) to generate instruction-enhanced CQA datasets. This engine comprises a data-filtering component (Sect. 5.1), utilizing a classifier with fine-grained chart features to reveal distributions and filter existing chart datasets. To mitigate the bias in the chart distributions and generate unavailable chart tasks, we further design a data generation component (Sect. 5.2) employing a chart space-guided data augmentation strategy to ensure the inclusivity of real-world possible charts. We further enrich reasoning-oriented QAs for the generated charts, contributing to a new CQA dataset and benchmark (Sect. 5.3) that features a wider variety of chart types and more QAs with effective visual instructions.
Existing MLLMs, mostly relying on CLIP encoders trained on natural images, are not optimally suited for visualization charts due to inherent differences. Recognizing the limitations, we develop a new MLLM (Sect. 6) that unfreezes the vision encoders in CLIP to better adapt to chart-specific features. Our MLLM is trained using the newly curated CQA dataset with more effective visual instructions. Additionally, we incorporate a mixture-of-resolution adaptation strategy[52] to enhance the fine-grained recognition capabilities of chart elements. Quantitative experiments (Sect. 7) demonstrate that even trained on a dataset with significantly less CQA data, our model consistently outperforms state-of-the-art CQA models on established benchmarks.
@@
  • An MLLM that outperforms existing open-source CQA models on existing CQA benchmarks and is comparable to the best commercial models on our proposed benchmark.
@@ 3 Related Work
The advent of MLLMs has shifted the paradigm, achieving breakthroughs in visual question answering[4]. Notably, the open-source generic model Qwen-VL[5] demonstrates superior performance over all specialized chart models in the ChartQA benchmark[54], especially those posed by humans as opposed to machine-generated questions. Despite these advancements, our extensive empirical study has uncovered limitations in the current MLLMs’ ability to handle real-world CQA tasks, especially those that fall outside of the training data distribution. Rectifying these limitations necessitates the consideration of the visualization reference model[9] when constructing training data, which elucidates the practical mapping process from raw data to final graphical representations. Accordingly, this study contributes to enhancing the performance of MLLMs in CQA by integrating knowledge from the visualization reference process into training data generation and augmentation.
Enhancing Capabilities of MLLMs. The enhancement of MLLMs in specific scenarios, such as medicine images and text-dense images, can be categorized into two primary approaches: model-centric works that aim to improve the performance and efficiency of vision encoders or projectors; data-centric works that try to improve the model performance by boosting the number and quality of training data. In data-centric advancements, several studies employ powerful LLMs (e.g., GPT-4[3]) to generate various instruction-format VL tasks, like caption generation[48]. Another line of studies has explored converting classical VL task datasets (e.g., COCO[42]) into an instruction-following format with pre-defined templates. Within this context, to enhance chart comprehension, ChartLLaMA[24] finetunes LLaVA with 160K instruction data generated by GPT-4. Similarly, ChartAst[57] crawls a huge amount of tables from arXiv and then uses tables to generate charts for large-scale chart-to-table pre-training. ChartAst also generates QA pairs based on the tables they collected. Despite these efforts, the factors contributing to efficient instruction data for chart understanding are still unclear.
@@
In the visualization field, real-world image datasets like Beagle[6], VisImage[18], Vis30K[12], multi-view [14] and composite visualizations [16, 22], and dashboards [17, 65], together with practical QA benchmarks for visual literacy test[36, 60, 21], have been introduced. The challenge lies in converting them into high-quality instruction data due to sparse label annotations. Our research draws upon methodologies that utilize GPT to generate code-format charts and associated instruction data. Specifically, we aim to guide the data generation process with the well-defined chart-task space[36] to contribute a dataset encompassing the real-world spectrum of chart features and QA tasks, thereby improving current MLLMs’ chart understanding capability.
@@ 4.1.2 Impacts of Distribution Bias
To gain deeper insights into the underlying reasons for the issue, we examine the responses generated by MLLMs equipped with CoT. Figure 3 illustrates three typical cases, highlighting deficiencies in three categories: recognition errors, inference errors for numerical calculations, and inference errors regarding chart knowledge. Multiple factors contribute to these errors. First, errors often occur for chart types common in visual literacy but rare in ChartQA, such as stacked bar charts. Additionally, uncommon questions in ChartQA, such as accurately determining a range of data values, may lead MLLMs to struggle to identify the correct range.
@@ 4.2.2 Results and Analysis
Table 2 shows the results of the ablation experiment of baseline and the fine-tuned MLLMs on different question types in ChartQA-H test set. Overall, models fine-tuned with more training data (individual vs. combinations datasets) achieve higher accuracy. Specifically, the inclusion of the human-generated ChartQA-H dataset substantially enhances model performance across all question types. In contrast, ChartQA-M dataset is less effective and mainly improves data retrieval and visual questions. This difference further underscores the limited impact of data retrieval questions for tackling CQA challenges and the critical role of diverse, reasoning-intensive questions over simple recognition questions. Moreover, Chart2Table serves as an accompanying effective instruction task if the data tables are available.
@@ 5 Data Engine
Total existing 611,655 5,334,247
Filtered dataset 69,418 68,223
@@ 5.1 Data filtering
Given that studies have shown common pre-trained visual feature extractors (e.g., CLIP-Vit[61]) are not sensitive to fine-grained chart attributes[72], conventional filtering approaches that sample data in the pre-trained feature space lead to inhomogeneity in these attributes. Additionally, most existing datasets lack detailed annotations beyond coarse-grained chart types (e.g., bar, line, and pie), posing challenges for stratified sampling. To mitigate this issue, we construct classifiers to learn those attributes in a supervised manner and then perform stratified sampling based on the labels predicted by the classifiers.
@@ 5.1.2 Image Sampling and Instruction Data Sampling
Figure 4 [Uncaptioned image] illustrates the image sampling process. We employ the CLIP-Vit[61] and a color extractor [1] to extract the overall feature and color feature of each image and then concatenate the two feature vectors to formulate a joint embedding space. Inspired by Bunny[26] and SemDeDup[2], we cluster images into $k$ clusters via $k$-means within the joint embedding space, aiming to group charts with similar features. To ensure chart attribute balancing, we incorporate stratified sampling within each cluster. Specifically, we create strata within each cluster according to predicted chart attributes and further perform sampling in each stratum. We identify duplicates by constructing an undirected graph, where edges connect image pairs with cosine similarity above a specified threshold $\mathcal{\varepsilon}$, indicating high feature similarity. We streamline the process by directly retaining only the image with the lowest cosine similarity to the stratum’s centroid from each set of semantic duplicates, thereby effectively reducing dataset size while preserving diversity. Finally, we manually adjust $\mathcal{\varepsilon}$ to obtain approximately 69K charts, ensuring an appropriate training cost.
@@ 5.3 Visualization-referenced Dataset and Benchmark
Overview of Dataset. Our generated dataset comprises 11 chart types and 8 task categories, as outlined in visual literacy research [36, 21, 60]. Table 4 illustrates the chart-task space. The generated dataset includes 10,385 table-chart pairs and 51,245 chart-QA pairs. By integrating the generated dataset with the filtered dataset (shown in Table LABEL:tab:existing-datasets), we ultimately produce a dataset of 199K, which includes 80K table-chart pairs and 119K chart-QA pairs. In contrast to earlier datasets that relied on single sources or template-driven designs, our approach effectively combines our generated dataset with existing datasets derived from a wide range of real-world and synthetic sources. This integration features a diverse array of visual encodings, referencing real-world charts from multiple sources. We have made deliberate efforts to ensure an equitable distribution of chart types and QA pairs, with a particular emphasis on underrepresented chart types and question types in current datasets, such as range determination and distribution characterization, to provide a more comprehensive and balanced resource. Detailed chart and QA examples can be found in Supplementary Section S2.
@@ 6.1 Model Architecture
High-resolution input by mixing vision encoders with tokens compression. The employed two-layer MLP effectively connects the feature space of vision encoders and LLMs but results in visual tokens that positively correlate with the image resolution. For instance, CLIP-Vit-L-14 results in 5,329 tokens for a 1,022×1,022 resolution image, as each token corresponds to a 14×14 image patch[52], which is computationally expensive for MLLMs. The popular query-based strategy for implementing high-resolution input, QFormer[40], requires large-scale pre-training to achieve vision-language alignments, which is impractical for visualization scenarios as high-quality data is scarce. Therefore, we adopt a resolution-adaptation strategy[52] to improve the resolution while supporting training on normal scale data. The strategy embeds high-resolution features into the low-resolution features via adapters, thus reducing visual feature tokens. Specifically, following the settings of LLaVA-HR[52], we integrate CLIP-ViT-L[61] and CLIP-ConvNeXt[50] as a mixture of vision encoders and then mix their features with an adaptation strategy, thus maintaining control over the length of the visual token sequence. The resolutions of ViT and CNN are set to 448×448 and 1,024×1,024, respectively.
@@ 6.2 Training Settings
Unfreezing the vision encoders. Former works[48, 24, 45] choose to freeze the vision encoder in the whole training process as the pre-trained CLIP is already good at capturing features of natural images. Their MLLM training target is aligning the features extracted by CLIP to the LLM embedding space by tuning the projector and LLM. However, former research[72] has found that CLIP performs much worse in visualization images, as its pre-training corpus has a relatively small amount of charts with coarse annotation, leading to limited chart recognition ability without further tuning. Unfreezing CLIP’s parameters enables better adaptation to chart features and improves the overall performance of MLLM’s chart understanding for improved chart recognition ability.
Training data. We skip the pre-training process and directly leverage the initial projector weight of LLaVA-HR[52] to conduct instruction tuning. Our study aims to improve general MLLM’s chart comprehension ability. Therefore, our training data consists of two parts: the 665K original instruction tuning data of LLaVA-1.5 and the 199K chart-related data described in Sect. 4.
Hyperparameters settings. AdamW[33] is used as the optimizer, and the learning rate (LR) and the global batch size are set to 2e-5 and 128, respectively. The training epoch is set to 1. The LR scheduler is cosine decay, with a warmup ratio of 0.03. The training is running on 16×NVIDIA A800 for approximately 19 hours.
@@ 7.1 Chart Attributes Classification
Table 6: Results on our benchmark. Bold and underlined numbers indicate the highest and second-highest metric values, respectively.
@@
No Filtered Data 91.5 60.9 76.2 90.9
No Generation Data 92.6 62.7 77.65 91.2
Our model (199K CQA data) 93.6 63.6 78.6 91.8
@@ 7.2.1 Benchmarks
ChartQA. The dataset information, the relaxed accuracy metric, and the prompt for short answers have been illustrated in Sect. 4.1.2. ChartQA’s training set is included in our training data.
@@ 7.2.3 Results
Table 7 presents the results of our model’s performance against other models. It demonstrates that our model consistently outperforms the baseline across all tasks. Particularly, we surpass the current leading models while utilizing significantly less data, showcasing our data filtering and generation effectiveness. Moreover, the ablation studies present the effectiveness of unfreezing vision encoders and mixing vision encoders for high-resolution.
@@ 8 Discussion
Visual encoder enhancement. Our research finds that unfreezing the vision encoders substantially enhances the chart recognition capabilities of MLLMs, showing the original CLIP-Vit’s under-performance in chart images[72]. An intuitive alternative design is replacing CLIP-Vit with an encoder pre-trained specifically on chart images. For example, ChartInstruct-LLaMA [56] substitute the CLIP-Vit in LLaVA with the UniChart vision encoder[55]. However, the researchers did not observe model performance improvements compared to Unichart. This highlights the superiority of generic vision encoders, which learn robust image interpreting capabilities (e.g., localization) from millions of natural images. Moreover, understanding some real-world charts requires broad vision knowledge. For instance, infographics may incorporate natural images to depict certain chart elements[73] vividly. Borrowing LLaVA-Med’s lesson[38] in initializing CLIP-Vit, developing a visualization-domain CLIP with enhanced basic chart interpretation performance is a promising future work.
@@
Insights on applying MLLMs to complex reasoning visualization tasks. Our research finds that current MLLMs still face challenges in analytical tasks (e.g., find anomalies and determine range). Recently, referential question-answering [76, 13] has been shown to benefit the comprehension of complex spatial relationships. It requires annotating bounding boxes and arrows in images and referring to these elements in questions. This task was not considered in our training data because of the lack of chart data with referential annotations. However, referential QAs are common in real-world visual analytics and potentially beneficial for tasks like find anomalies. For example, we can use bounding boxes to label anomaly points or highlight a range of data elements, boosting MLLMs’ understanding of relevant tasks. We place the exploration of referential and other possible complex QA formats tailored for charts in future work. Furthermore, end-to-end MLLM outputs inherently pose uncertainty. Incorporating golden tables and code of charts with MLLM for interaction applications may be robust in complex visual analytics scenarios.